luhivqqcherry:

more mathh workk last questionn . Write in summation notation: 3 + 6 + 9 + 12 + ... + 30 a) C b) A c) B

luhivqqcherry:

1 attachment
Midnight97:

Alright, working on it

Midnight97:

Let me work it out rq and I'll get it to you (:

luhivqqcherry:

@midnight97 wrote:
Let me work it out rq and I'll get it to you (:
yess okayy !! .

Midnight97:

so summation notation, often represented with the sigma symbol Σ , is a way of expressing the sum of a series of terms. The series we have is an arithmetic sequence where each term increases by a certain number, which in our case is 3 To write your series in summation notation, we'll need to identify: 1. The first term... which is 3

Midnight97:

So that's what we'll call a1 The next thing we need to identify is the numerical difference between each number... which is d=3

Midnight97:

Because if you look at your question the numbers are going up by 3

luhivqqcherry:

@midnight97 wrote:
Because if you look at your question the numbers are going up by 3
yesss .

Midnight97:

Now it's said that there are "n" terms in the sequence and the last thing we need to know and fill in is the general sequence formula

Midnight97:

luhivqqcherry:

this makes my brain hurt .

luhivqqcherry:

but so that means it's A ? .

Midnight97:

Now we have to figure out what n equals Since the last term is 30 and we start at 3 and increase by 3 each time, we can find n by solving the equation for 30 so: 30=3+(n-1)x 3 once a1 and d are plugged in and now we solve 30= 3+3n -3

Midnight97:

30= 3+3n -3 becomes 30=3n So n= 1

Midnight97:

Wait wait I did that wrong

Midnight97:

@midnight97 wrote:
30= 3+3n -3 becomes 30=3n So n= 1
n= 10

luhivqqcherry:

@midnight97 wrote:
@midnight97 wrote:
30= 3+3n -3 becomes 30=3n So n= 1
n= 10
OHHHHHH OKAYYY .

Midnight97:

Your answer is B

luhivqqcherry:

YAYYAYAYYYY TYYTYTYSM FOR ALLL THE HELPP .

Midnight97:

@luhivqqcherry wrote:
YAYYAYAYYYY TYYTYTYSM FOR ALLL THE HELPP .
You're welcome